OpenStudy (anonymous):

7.2-7). Divide the following expressions. Be sure to reduce all answers to lowest terms. 6y^2+21y+9 12y^2+24y+9 ____________ divide by _____________ y^2-16 2y^2-5y-12

OpenStudy (anonymous):

each numerator and denominator can be factored

OpenStudy (anonymous):

6*y^2+21y+9=3*(2y+1)(y+3) y^2-16=(y+4)(y-4) 12y^2+24y+9=3*(2y+3)(2y+1) 2y^2-5y-12=(2y+3)(y-4) putting it all together we see that the 3 in front, (2y+3), (2y+1), and (y-4) all cancel. We are left with (y+3)/(y+4)